OpenStudy (anonymous):

In 2004 there were 72 million voters registered as Democrats, 55 million American voters registered as Republicans, and 42 million registered Independents. What is the probability that 3 randomly selected, registered voters would be Independents?

OpenStudy (kropot72):

The total number of voters is (72 + 55 + 42) million. The probability of a randomly selected voter being an independent is \[P(Independent)=\frac{42}{72+55+42}=you\ can\ calculate\] When you have posted the result of the above calculation we can move on to the final step in the solution.

OpenStudy (anonymous):

i got .24852071

OpenStudy (anonymous):

0.15 0.077 0.015 0.036 but idk how when these are the ones i have to choose from its frustrating

OpenStudy (kropot72):

Correct! When selecting each of the voters the probability that they are an Independent is 0.24852071. The result of each selection is independent of the other results. Therefore the probability of all three being Independents is given by \[P(all\ 3\ Independents)=(0.248520710)^{3}=you\ can\ calculate\]

OpenStudy (anonymous):

so would i do 169 by what i got

OpenStudy (kropot72):

You just need to calculate 0.24852071 * 0.24852071 * 0.24852071 = ?

OpenStudy (anonymous):

so it would be .015 right

OpenStudy (kropot72):

Good work! Your result is correct.

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)